(Please show calculations for study purposes)

Bandar Industries Berhad of Malaysia manufactures sporting equipment. One of the company’s products, a football helmet for the North American market, requires a special plastic. During the quarter ending June 30, the company manufactured 3,300 helmets, using 2,244 kilograms of plastic. The plastic cost the company $17,054.

According to the standard cost card, each helmet should require 0.63 kilograms of plastic, at a cost of $8.00 per kilogram.

Required:

1. What is the standard quantity of kilograms of plastic (SQ) that is allowed to make 3,300 helmets?

2. What is the standard materials cost allowed (SQ × SP) to make 3,300 helmets?

3. What is the materials spending variance?

4. What is the materials price variance and the materials quantity variance?
Answer #1

Standard Quantity allowed = 3300*0.63

= 2079

Standard cost allowed = 2079*8

= 16632

Materials spending variance = 16632-17054

= 422 U

Materials Price Variance = (SP-AP) *AQ = (8-17054/2244)*2244

= 898 F

Materials Quantity Variance = (SQ-AQ) *SP= (2079-2244)*8

= 1320 U